OpenStudy (avanti):

Evaluate tan(Sin-1(-5/13)). Answer as a fraction please.

OpenStudy (zzr0ck3r):

Sorry on tablet and the site won't let me draw

OpenStudy (kinggeorge):

There should be a particular triangle that has two sides of length 5 and 13. Can you tell me what the third side length is?

OpenStudy (avanti):

12 is the third side

OpenStudy (kinggeorge):

Right. So you can draw a triangle with side lengths 5,12,13. And \(\sin(x)=5/13\). Now can you tell me what \(\tan(x)\) is?

OpenStudy (avanti):

tan(x)=5/12

OpenStudy (kinggeorge):

Bingo. Now we just need to fit these pieces together.\[\sin(x)=5/13\implies\sin^{-1}(5/13)=x\implies\sin^{-1}(-5/13)=-x\]Then we take the tangent.\[\tan(\sin^{-1}(-5/13))=\tan(-x)=-5/12\]All this make sense?

OpenStudy (avanti):

oh okay thank you! how come you made the 5/13 and the x negative in the third part of the 2nd row?

OpenStudy (kinggeorge):

If you look at your original problem, we want to find \(\tan(\sin^{-1}(-5/13))\), which has a negative in front of the 5/13. So I made the 5/13 negative. Then, \(\sin\) has the property where \(\sin(-x)=-\sin(x)\). So I applied that to get the negative in front of the x.

OpenStudy (avanti):

okay that makes sense, thank you very much!

OpenStudy (kinggeorge):

You're very welcome.